OpenStudy (anonymous):

The Nut Barn requires 15 pounds of mixed walnuts and pecans from their wholesaler. Walnuts cost $2.26 per pound and pecans cost $1.96 per pound. How many pounds of pecans are needed if they only want to pay $2.18 per pound for the mixture?

OpenStudy (anonymous):

(1) W +P = 15 (2) (2.26W+1.96P)/15=2.18 ------------------------- (2) 2.26 W+1.96 P=2.18*15 (2) 2.26 W +1.96 P =32.7 (2) 226 W + 196 P = 3270 Multiply both sides of (1) by 196 and subtract (1) from (2) (2) 226 W +196 P =3270 (1) -196 W - 196 P =-2940 30 W = 330 W = 11 and, since (1) W + P = 15 (1) P = 4 11 pounds of walnuts and 4 pounds of pecans

OpenStudy (anonymous):

OH THANK YOU SO MUCH!!